OpenStudy (anonymous):

The table shows the amount of green paint that can be added to different amounts of white paint while making a custom paint color: Paint Color Quarts of Green Paint Quarts of White Paint 3 6 4 8 5 10 6 12 Which statement is true about the ratio of green paint to white paint in the paint color? It is 1:2. It is 2:1. It is 1:3. It is 3:1.

OpenStudy (anonymous):

If you want to find a ratio between two numbers you have to take the one pair of values (e.g. 6 and 8) and set them equal to another pair of values (e.g. 3 and 4) as fractions. The values used have to be pairs and cannot be any random numbers. For example, 6/8=3/4. To solve this ratio, simply cross multiply and you have your answer!

OpenStudy (anonymous):

Also if you want to check to see if your ratio is constant throughout the data set, set the first and last pairs up as fractions and solve, so 3/6=6/12. This gives you your average rate of change and makes sure that the ratio is constant throughout.